# Understanding Pre-training and Fine-tuning from Loss Landscape Perspective

### v1

@@ 4.2 Input-space Robustness Can Be Bounded by Average-case Alignment
Let $\mathcal{D}^{\prime}$ be a modified version of $\mathcal{D}$ where $k$ tokens are substituted, i.e., each token $\bm{e}_{i}$ is replaced with $\bm{e}_{i}^{\prime}$ in the set $\mathcal{C}=\{(\bm{e}_{i},\bm{e}_{i}^{\prime})\}_{i=1}^{k}$. For any benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, with probability at least $1-\delta$, the performance degradation can be bounded by the embedding differences:
$J_{\mathcal{D}^{\prime}}(f_{\bm{\theta}+\bm{\epsilon}})\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})]\right)-\frac{\sqrt{\sum_{i=1}^{k}\|\bm{W}\bm{e}_{i}-\bm{W}\bm{e}_{i}^{\prime}\|_{2}^{2}}}{\sigma}\right)-L\sigma\sqrt{2\log\frac{1}{\delta}}.$ (10)
A straightforward application of this theorem involves comparing the modified and original inputs, calculating the equivalent weight differences, and applying the results from Sec. 4.1.
Application to Certification Against Jailbreaking. Consider $\mathcal{D}=\{\bm{x}\}$ containing a single input instance and $J$ as a safety detector that returns values greater than 0.5 for safe outputs and less than 0.5 for harmful outputs. We can certify robustness against jailbreaking attacks by determining whether $J_{\mathcal{D}^{\prime}}(f_{\bm{\theta}+\bm{\epsilon}})$ remains above 0.5 after modifying $\mathcal{D}=\{\bm{x}\}$ to $\mathcal{D}^{\prime}=\{\bm{x}_{adv}\}$.
We derive the following conclusions for Qwen2.5-7B [71]:
Substituting Special Tokens Preserves Performance. In Qwen2.5-7B [71], special tokens such as "<s>", ".", "···", " ", and " " can be interchanged without compromising performance. Due to their small L2 distances, substituting these tokens, as evaluated in Eq. 10, results in equivalent parameter perturbations too small to significantly alter model outputs. Thus, the model exhibits robustness to these special or infrequently used tokens.
Substituting Tokens With or Without Prefix Spaces Has Limited Impact. Most large language models, including Qwen2.5-7B, use BPE tokenizers [61], where tokens with and without leading spaces are distinct (e.g., "hi" vs. " hi"). During decoding, the tokenizer automatically adjusts based on sentence position. Many certifiable substitution pairs include tokens with and without leading spaces, indicating minimal performance impact.
Current LLMs Are Generally Sensitive to Input Changes. As shown in Fig. 4(c), only a small fraction of token substitutions have negligible effects on model outputs. This flexibility allows models to generate diverse responses based on input variations but also introduces robustness vulnerabilities. Minor input changes can lead to significant output variations, potentially resulting in incorrect or unsafe responses [77, 14, 3].
@@ 4.3 Basin May Have Sufficient Expressive Power in the Future
Given the theoretical guarantees of fine-tuning within a basin, one might wonder whether subsequent fine-tuning can be constrained to a theoretically guaranteed region, thereby preserving all original capabilities and achieving continual learning without forgetting [15].
A primary concern with this constraint method is whether it limits the expressive power of the hypothesis set. Specifically, if the hypothesis set is constrained to functions where $\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}\leq O(\sigma)$, can it still effectively fit the fine-tuning dataset?
We propose that a large model with a parameter norm constraint of $O(\sigma)$ may exhibit greater expressive power than smaller models with significantly larger norms. This is because expressive power primarily depends on the number of parameters or layers, rather than the norm of the parameters. Thus, training a larger model could address concerns about expressive power.
@@
([7]) The upper and lower bounds of the VC dimension of an $l$-layer fully connected ReLU network are $O(dl\log d)$ and $\Omega(dl\log\frac{d}{l})$, respectively.
@@
(Adapted from [51]) The upper bound of the Rademacher complexity of a two-layer ReLU neural network, where each parameter has a value bound of $O(\sigma)$, is $O(d\sigma^{2})$.
5 Enlarging the Alignment Basin
Given that a larger basin provides stronger robustness guarantees against any fine-tuning and jailbreaking attacks, one might wonder whether the basin can be enlarged. In this section, we propose that the basin can be readily expanded, may due to the over-parameterization property of neural networks [43, 2].
@@ 5.1 Gaussian-augmented Optimizer
As outlined in Definition 4, enlarging the basin size requires optimizing $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{\mathcal{D}}(f_{\bm{\theta}+\bm{\epsilon}})]$, ensuring that the model $\bm{\theta}$ is robust to Gaussian perturbations. To this end, we define the loss function as the expected cross-entropy loss over perturbed parameters:
@@
As detailed in Algorithm 1, the optimization process involves performing a forward pass with perturbed parameters $\bm{\theta}+\bm{\epsilon}$, computing the loss, calculating the gradient via backpropagation, and using the gradient to update the parameters with a standard optimizer.
Algorithm 1 Gaussian-augmented Optimizer for Basin Enlargement (GO optimizer)
0:  Model $f_{\bm{\theta}}$, dataset $\mathcal{D}$, perturbation variance $\sigma^{2}$, base optimizer (e.g., SGD or Adam)
1:  for each gradient step do
2:     Sample mini-batch $\{\bm{x}_{i}\}_{i=1}^{B}\sim\mathcal{D}$ and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$.
3:     Compute gradient on perturbed parameters $\nabla_{\bm{\theta}}L_{\text{train}}=-\sum_{i=1}^{B}\nabla_{\bm{\theta}}\log p(\bm{x}_{i}|\bm{\theta}+\bm{\epsilon})$
4:     Update parameters $\bm{\theta}\leftarrow\text{Optimizer}(\bm{\theta},\nabla_{\bm{\theta}}L_{\text{train}})$
5:  end for
6:  return  $\bm{\theta}$
@@ 5.2 Experimental Results
Settings. We utilize the UltraFeedback dataset [19], which comprises diverse corpora designed to endow models with foundational knowledge and conversational capabilities. Due to limited computational resources, we fine-tune a compact model, Qwen2.5-0.5B, using the Gaussian-augmented optimizer outlined in Algorithm 1 for four epochs.
Figure 5: The loss landscape of original Qwen-2.5-0.5B and fine-tuned by GO optimizer with $\sigma=0.005$ and $\sigma=0.01$.
Results. As shown in Fig. 5, the original Qwen2.5-0.5B model fails to maintain conversational abilities when its parameters $\bm{\theta}$ are perturbed by Gaussian noise with a variance of 0.05, producing only incoherent outputs. In contrast, when trained using the Gaussian-augmented optimizer with $\sigma=0.005$ or $\sigma=0.01$, the model sustains basic conversational capabilities under the same noise perturbation. However, these models perform similarly to a model trained directly on the UltraFeedback dataset without pre-training; that is, using the Gaussian-augmented optimizer only during the fine-tuning phase, rather than both pre-training and fine-tuning, does not enable the perturbed model $f_{\bm{\theta}+\bm{\epsilon}}$ to perform comparably to the unperturbed model $f_{\bm{\theta}}$, since these parameters $\bm{\theta}+\bm{\epsilon}$ have not benefit from pre-training. This suggests that the Gaussian-augmented optimizer is most effective during the pre-training phase, enabling the model to retain capabilities across a broader parameter neighborhood. See Appendix G for further details.
@@ 6 Conclusion
In this work, we explore the loss landscape of large language models to elucidate the alignment brittleness phenomenon. We demonstrate that the loss landscape of large language models resembles a basin, within which models perform nearly identically and outside of which they lose all capabilities. This property enables us to derive a theoretical lower bound on performance degradation during any fine-tuning and jailbreaking attacks within certain norm constraints. We also show that the basin can be readily expanded due to the over-parameterization property of neural networks. Despite these contributions, our exploration remains preliminary. We hope our work sheds light on the alignment brittleness phenomenon and motivates large-scale studies that apply the Gaussian-augmented optimizer during the pre-training phase of larger models, verifying the practical utility of theoretical guarantees, the practical relationship between the robustness of parameters against Gaussian perturbations and general fine-tuning.
@@ Appendix A Notations
$b$ Auxiliary variable that only for finding worst-case direction.
@@
$\Phi^{-1}$ CDF of the standard Gaussian distribution.
@@ Appendix B Related Work
This work is greatly inspired by [54], which argues that the safety loss landscape resembles a basin, within which the model is safe and outside of which it is not. Our study extends [54] by investigating the loss landscape across additional capabilities and providing a deeper analysis of the relationship between loss landscapes and catastrophic forgetting during fine-tuning.
Concurrent work [63] suggests that over-pre-trained large language models are harder to fine-tune because they lack robustness to parameter perturbations. Our work complements this study by explaining how robustness to Gaussian parameter perturbations provides a lower bound on performance degradation during fine-tuning.
@@ Appendix C Why Do the Loss Landscapes of LLMs Resemble Basins?
This paper presents a seemingly different conclusion from [54], where we argue that the loss landscape of large language models resembles a basin, whereas they suggest that LLM performance degrades gradually as the perturbation budget increases. The key reason is that [54] uses benchmarks that evaluate models by log-likelihood, where the loss landscape is expected to be continuous and smooth. In contrast, we use benchmarks that evaluate models based on their generation results.
This raises the question of why large language models perform nearly identically within a certain range of Gaussian perturbations to their parameters but rapidly lose all capabilities when perturbations exceed this range. This phenomenon aligns with recent findings that large language models can resist common noise perturbations [49, 29, 64], which show that removing certain modules in LLMs leaves their capabilities largely intact, but removing more causes a swift loss of conversational ability.
In this work, we do not have a definitive answer to this question. However, we hypothesize that this may stem from the over-parameterization property of LLMs. Mode connectivity is an important property of over-parameterized models, stating that the modes of an over-parameterized network are connected to one another [26, 25, 45]. For large language models, beyond permutation invariance in feed-forward networks [1], there is also rotational invariance in key and query matrices. Consequently, these equivalent modes may be closer to each other, forming a subspace with dimensions nearly equivalent to the original space. We leave this exploration to future work.
@@ D.1 More Experimental Details
Normalizing the Loss Landscape. As clarified in Sec. 3.1, since the values of different benchmarks are not directly comparable, we normalize each landscape to the interval $[0,1]$ and invert benchmarks where higher values indicate better performance, unifying them such that lower values indicate better performance for consistent visualization. This normalization involves three steps. First, we subtract the minimum value. Then, we divide by the maximum value. Finally, we invert the value.
D.2 Loss Landscapes of Large Models
@@
Figure 6: The most-case loss landscape of three Qwen2.5 models with different sizes. The larger the model, the larger the basins.
@@
We also visualize the loss landscapes of larger models, as shown in Fig. 6. We draw the following conclusions:
@@
Larger models tend to have larger basins. As shown, for each capability, including basic, math, safety, and coding, the basic capability basin of Qwen2.5-0.5B is small, while that of Qwen2.5-7B is larger. Qwen2.5-32B has the largest basin, nearly twice the size of Qwen2.5-7B.
@@
Larger models have greater expressive power within their basins. As analyzed in Sec. 4.3, the larger the model and its basin, the greater the expressive power. Since larger models have both more parameters and larger basin sizes, they exhibit significantly greater expressive power than smaller models.

### v2

@@ 4.2 Input-space Robustness Can Be Bounded by Most-case Alignment
Let $\mathcal{D}^{\prime}$ be a modified version of $\mathcal{D}$ where $k$ tokens are substituted, i.e., each token $\bm{e}_{i}$ is replaced with $\bm{e}_{i}^{\prime}$ in the set $\mathcal{C}=\{(\bm{e}_{i},\bm{e}_{i}^{\prime})\}_{i=1}^{k}$. For any benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, the performance degradation can be bounded by the embedding differences:
$\mathbb{E}_{\bm{\epsilon}}[J_{f,\mathcal{D}^{\prime}}(\bm{\theta}+\bm{\epsilon})]\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)-\frac{\sqrt{\sum_{i=1}^{k}\|\bm{W}\bm{e}_{i}-\bm{W}\bm{e}_{i}^{\prime}\|_{2}^{2}}}{\sigma}\right).$ (7)
A straightforward application of this theorem involves comparing the modified and original inputs, calculating the equivalent weight differences, and applying the results from Sec.˜4.1.
Application to Certification Against Jailbreaking. Consider $\mathcal{D}=\{\bm{x}\}$ containing a single input instance and $J$ as a safety detector that returns values greater than 0.5 for safe outputs and less than 0.5 for harmful outputs. We can certify robustness against jailbreaking attacks by determining whether $J_{f,\mathcal{D}^{\prime}}(\bm{\theta}+\bm{\epsilon})$ remains above 0.5 after modifying $\mathcal{D}=\{\bm{x}\}$ to $\mathcal{D}^{\prime}=\{\bm{x}_{adv}\}$.
We derive the following conclusions for Qwen2.5-7B (Yang et al., 2024):
Substituting Some Tokens Preserves Performance. Most large language models, including Qwen2.5-7B, use BPE tokenizers (Sennrich, 2015), where tokens with and without leading spaces are distinct (e.g., “hi” vs. “_hi”). These tokens often have small $\ell_{2}$ distances between each other. Special tokens, such as “<s>”, “.”, “”, “ ”, and “ ”, also have small $\ell_{2}$ distances. Substituting these tokens, as evaluated in Eq.˜7, results in equivalent parameter perturbations too small to significantly alter model outputs. Thus, the model exhibits robustness to these token pairs.
Current LLMs Are Generally Sensitive to Input Changes. As shown in Fig.˜4(c), only a small fraction of token substitutions have negligible effects on model outputs. This flexibility allows models to generate diverse responses based on input variations but also introduces robustness vulnerabilities. Minor input changes can lead to significant output variations, potentially resulting in incorrect or unsafe responses (Zou et al., 2023; Chen et al., 2025a; Andriushchenko et al., 2024).
(a) Pretraining Loss Curve
(b) Loss Landscape
(c) SFT Loss Curve
Figure 5: Pre-training and fine-tuning using GO and Adam optimizers. (a) The pre-training loss curve on OpenWebText. (b) The loss landscape on OpenWebText after pre-training. (c) Performance changes in old capability (OpenWebText) and new capability (Alpaca).
@@ 4.3 Basin May Have Sufficient Expressive Power in the Future
Given the theoretical guarantees of fine-tuning within a basin, one might wonder whether subsequent fine-tuning can be constrained to a theoretically guaranteed region, thereby preserving all original capabilities and achieving continual learning without forgetting (Chen et al., 2025b). A primary concern with this constraint method is whether it limits the expressive power of the hypothesis set. Specifically, if the hypothesis set is constrained to functions where $\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}\leq O(\sigma)$, can it still effectively fit the fine-tuning dataset?
We propose that the larger the model and its basin, the greater the expressive power within the basin, enabling better acquisition of new capabilities. As demonstrated in the following lemmas, a large model with a large basin may be even more expressive than a smaller model. Thus, in the future, one may prefer training a significantly larger model and enlarging its basin, either actively or passively (as larger models naturally form larger basins). In this case, constraining fine-tuning within the basin will provide sufficient expressive power to acquire new capabilities.
@@
(Bartlett et al., 2019) The upper and lower bounds of the VC dimension of an $l$-layer fully connected ReLU network are $O(dl\log d)$ and $\Omega(dl\log\frac{d}{l})$, respectively.
@@
(Adapted from Neyshabur et al. (2015)) The upper bound of the Rademacher complexity of a two-layer ReLU neural network, where each parameter has an absolute value bound of $O(\sigma)$, is $O(d\sigma^{2})$.
5 Enlarging Basins
Given that a larger basin provides stronger robustness guarantees against any fine-tuning and jailbreaking attacks, one might wonder whether the basin can be enlarged. In this section, we discover that the basin can be readily expanded, may due to the over-parameterization property of current LLMs (Liu et al., 2022; Allen-Zhu et al., 2019).
@@ 5.1 Gaussian-augmented Optimizer
As outlined in Definition 3, enlarging the basin size requires optimizing $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$, ensuring that the model $\bm{\theta}$ is robust to Gaussian perturbations. To this end, we define the loss function as the expected cross-entropy loss over perturbed parameters:
@@
As detailed in Algorithm 1, the optimization process involves performing a forward pass with perturbed parameters $\bm{\theta}+\bm{\epsilon}$, computing the loss, calculating the gradient via backpropagation, and using the gradient to update the parameters with a standard optimizer. We name this process GO optimizer.
@@ 5.2 Experimental Results
Settings. To better study the effect of the GO optimizer, we applied it during pre-training. Following the NanoGPT pipeline (Karpathy, 2023), we pre-train a GPT2-127M model on OpenWebText for 600,000 steps using the same hyperparameters as in the repository. For the GO optimizer, we use identical hyperparameters, except for the additional hyperparameter $\sigma=0.01$. We then fine-tune these two models on the Alpaca dataset to enhance conversational capabilities while examining performance degradation, both using the Adam optimizer.
Pre-training Results. As shown in Fig.˜5(b), adding Gaussian noise to parameters can significantly expand the basin size. Although the GO optimizer is slower than Adam at the beginning of training, as it requires optimizing the loss across the entire neighborhood, it gradually catches up and reduces the performance gap (see Fig.˜5(a)), possibly due to the over-parameterization property of current large language models. More interestingly, the GO optimizer appears to introduce some inductive bias, outperforming AdamW on some benchmarks (see Sec.˜D.5).
Fine-tuning Results. We propose that the larger the basin, the less performance degradation occurs during benign fine-tuning (since benign fine-tuning, which aims to enhance the model, should not be more harmful than Gaussian noise) and that it does not affect the acquisition of new capabilities (since larger basins imply greater expressive power within the basin). As demonstrated in Fig.˜5(c), fine-tuning a model pre-trained with the Gaussian-augmented optimizer achieves significantly less performance degradation on OpenWebText while maintaining nearly the same learning speed on Alpaca, supporting this claim. See Sec.˜D.5 for detailed analysis of additional benchmarks.
@@ 6 Conclusion
In this work, we explore the loss landscape of large language models to elucidate the alignment brittleness phenomenon. We demonstrate that the loss landscape of large language models resembles a basin, within which models perform nearly identically and outside of which they lose all capabilities. This property enables us to derive a theoretical lower bound on performance degradation during any fine-tuning and jailbreaking attacks within certain norm constraints. We also show that the basin can be readily expanded due to the over-parameterization property of neural networks. Despite these contributions, our exploration remains preliminary. We hope our work sheds light on the alignment brittleness phenomenon and motivates large-scale studies that apply the GO optimizer during the pre-training phase of larger models at cutting-edge scales, verifying the practical utility of these theoretical guarantees and the relationship between basin size and performance degradation in general fine-tuning. See Appendix˜F for more detailed discussion.
@@ Appendix A Notations
$\Phi^{-1}$ Inverse CDF of the standard Gaussian distribution.
@@
$p_{\text{lower}}$, $p_{\text{upper}}$ Confidence interval obtained by Clopper Pearson Bound.
@@ Appendix B Related Work
This work is greatly inspired by Peng et al. (2024), which argues that the safety loss landscape resembles a basin, within which the model is safe and outside of which it is not. Our study extends Peng et al. (2024) by investigating the loss landscape across additional capabilities and providing a deeper analysis of the relationship between loss landscapes and catastrophic forgetting during fine-tuning.
Concurrent work (Springer et al., 2025) suggests that over-pre-trained large language models are harder to fine-tune because they lack robustness to parameter perturbations using Gaussian noise. Our work complements this study by explaining how robustness to Gaussian parameter perturbations provides a lower bound on performance degradation during fine-tuning.
(a) Basic
(b) Safety
(c) Math
(d) Coding
Figure 6: The 3D version of the most-case loss landscape, using Qwen2.5-7B model.
@@ Appendix C Why Do the Loss Landscapes of LLMs Resemble Basins?
This paper presents a seemingly different conclusion from Peng et al. (2024), where we argue that the loss landscape of large language models resembles a basin, whereas they suggest that LLM performance degrades gradually as the perturbation budget increases. The key reason is that Peng et al. (2024) uses benchmarks that evaluate models by log-likelihood, where the loss landscape is expected to be continuous and smooth (See Fig.˜8(a)). In contrast, we use benchmarks that evaluate models based on their generation results.
This raises the question of why large language models perform nearly identically within a certain range of Gaussian perturbations to their parameters but rapidly lose all capabilities when perturbations exceed this range. This phenomenon aligns with recent findings that large language models can resist common noise perturbations (Men et al., 2024; Gromov et al., 2024; Sun et al., 2023), which show that removing certain modules in LLMs leaves their capabilities largely intact, but removing more causes a swift loss of conversational ability.
In this work, we do not have a definitive answer to this question. However, we hypothesize that this may stem from the over-parameterization property of LLMs. Mode connectivity is an important property of over-parameterized models, stating that the modes of an over-parameterized network are connected to one another (Garipov et al., 2018; Frankle et al., 2020; Lubana et al., 2023). For large language models, beyond permutation invariance in feed-forward networks (Ainsworth et al., 2022), there is also rotational invariance in key and query matrices. Consequently, these equivalent modes may be closer to each other, forming a subspace with dimensions nearly equivalent to the original space. We leave this exploration to future work.
@@ D.1 More Experimental Details
Normalizing the Loss Landscape. As described in Sec.˜3.1, since benchmark values vary in scale and direction, we normalize each loss landscape to the interval $[0,1]$ and invert benchmarks where higher values indicate better performance, ensuring that lower values consistently represent better performance for unified visualization. This normalization involves three steps: first, subtracting the minimum value; then, dividing by the range to scale to $[0,1]$; and finally, inverting the value by computing one minus the normalized value.
D.2 Raw Numerical Values of Basins
$\alpha$ ($\times 10^{-3}$) 0 0.5 1 1.5 2 2.5 3 3.5 4 4.5 5 5.5 6 6.5 7
Safety 1 1 1 0.98 0.98 0.98 0.98 0.98 0.98 0.98 0.96 0.78 0.06 0 0
Math 0.84 0.84 0.84 0.84 0.81 0.7 0.72 0.58 0.32 0.21 0.13 0.01 0 0 0
Basic 0.76 0.76 0.76 0.76 0.73 0.72 0.67 0.65 0.60 0.53 0.43 0.32 0.25 0.16 0
Coding 0.88 0.89 0.89 0.86 0.86 0.83 0.8 0.74 0.74 0.54 0.45 0.11 0 0 0
Table 1: The raw benchmark value $J_{f,\mathcal{D}}(\bm{\theta}+\alpha\bm{\delta})$ ($\uparrow$, the higher the better) of the most-case landscape using Qwen2.5-7B. Due to page limitations, we present only half of the data here. Thus, the basin size is twice as large as shown, accounting for (nearly) symmetric counterpart on the other side. As shown, the loss landscape literally forms basins, within these basins, the benchmark values remain literally unchanged.
Settings. To better illustrate the basin phenomenon in the most-case loss landscape shown in Fig.˜1, we provide the raw benchmark values for the Qwen2.5-7B model. Due to page limitations, we present data only for Qwen2.5-7B and include less than half of the perturbation range for $\alpha$, since the full dataset exhibiting near-symmetric behavior on the other side.
Results. As shown in Table˜1, the most-case loss landscape of Qwen2.5-7B literally forms basins, where benchmark values remain completely unchanged within specific ranges of parameter perturbations. Specifically, for the Safety benchmark, the value remains constant at 1 for perturbations up to $\alpha=1\times 10^{-3}$, and for the Math and Basic benchmarks, the values are stable at 0.84 and 0.76, respectively, up to $\alpha=1.5\times 10^{-3}$. Similarly, the Coding benchmark shows near-constant values (0.88 to 0.89) up to $\alpha=1.5\times 10^{-3}$. This stability demonstrates that, within these basins, the model’s capabilities are entirely unaffected by random noise perturbations, aligning with our claim that the most-case landscape forms a robust basin structure, as visualized in Fig.˜1.
D.3 Loss Landscapes of Large Models
@@
Figure 7: The most-case loss landscape of six Qwen2.5 models with different sizes. As shown, for the 0.5B model, the landscape resembles a small model and does not even resemble a basin. When the models become larger, the basins also become larger and clearer.
@@
We also visualize the loss landscapes of larger models, as shown in Fig.˜7. We draw the following conclusions:
@@
Larger models tend to have larger basins. As shown, for each capability, including basic, math, safety, and coding, the basin of Qwen2.5-0.5B is small, while that of Qwen2.5-7B is larger. Qwen2.5-32B has the largest basin, nearly twice the size of Qwen2.5-7B.
@@
Larger models have greater expressive power within their basins. As analyzed in Sec.˜4.3, the larger the model and its basin, the greater the expressive power within the basin. Since larger models have both more parameters and larger basin sizes (may due to their over-parameterized property), they exhibit significantly greater expressive power than smaller models.